Calculate the algal mass present in one liter of your water sample using chlorophyll a value of 9.502ug/L (microgram/liter). First get the typical proportion of chlorophyll a found in an algal cell.
Answer #1

Algal biomass can be calculated using the content of chlorophyll. Typically in an algal cell chlorophyll a to chlorophyll b is 3:1.

In one liter water we have chlorophyll a- 9.502ug

So, chlorophyll b content is - 4.751ug

Chlorophyll content- a+b = 9.502+4.751 = 14.252

1 ug = 0.001 mg

1 liter = 0.001 cubic meter

conversion factor: 0.001/0.001 = 1

So, algal mass present in one litre of water = 14.252 cubic meter